A jet plane passes over you at a height of 4990 m and a speed of Mach 1.75. (a) Find the Mach cone angle. (b) How long after the jet passes directly overhead does the shock wave reach you? Use 331 m/s for the speed of sound.
Answer #1

a)speed of mach=1.75
let mach cone angle be a
then, sin(a)=1/1.75
thus a=34.85 degrees
b)let time after which shock wave reaches the receiver be t
then distance travelled by the plane=1.75*331*t metres
height=4990 m
then tan(a)=4990/(1.75*331*t)
=> tan(34.85)=4990/579.25t
=>t=4990/(403.34)s
=12.37 s